OpenStudy (vishweshshrimali5):

If x is the mole fraction of a solute when dissolved in pure waiter, then prove that the molality of solution is Molality (m) = (500x)/(9-9x) PLEASE HELP ITS URGENT

OpenStudy (anonymous):

i typed it into my fancy calculator an it said its 0 idk if its rite an idk how to do it

OpenStudy (trancenova):

Hi, have been trying to work it out but am a little stuck. Here is what I have so far (need to go to bed!) but I think I'm on the right track, probably has a few errors but hopefully you can do the rest :) m= n/v x = n/m x = number of sol molecules (n) / number of total molecules (n + w) M = number of sol molecules (n) / volume (v) M = (500x)/(9-9x) n/v = 500x/(9-9x) number of molecules in a volume of water = ? room temp (water) 1g = 1ml 1L = 1000mL water density = 1000g/L 1000g/L / 18g/mole = 55.555 using Avagadro 55.555 * (6.022*10^23) = 334.55 * 10^23 molecules in 1L of water x = n/(n+ 3.34*10^25) xn + 3.34*10^25x = n 3.34*10^25x= n - xn 3.34*10^25x = n(1-x) n = 3.34*10^25x/(1-x) M= n/v M = (3.34*10^25x/(1-x))/V now as we were using V = 1L 3.34*10^25x/(1-x) = (500x)/(9-9x)